OpenStudy (anonymous):

Use the limit definition of the derivative f(x)=1/(x+3) f'(-6) f'(-4)

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

please help me :(

zepdrix (zepdrix):

\[\large \color{orangered}{f(x)=\frac{1}{x+3}}\]The Limit Definition of the Derivative:\[\large f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-\color{orangered}{f(x)}}{h}\] See the part that I colored orange in the definition? We can plug that right in. But we also need to plug in \(f(x+h)\) correctly.

zepdrix (zepdrix):

\[\large f(\color{royalblue}{x})=\frac{1}{\color{royalblue}{x}+3}\]\[\large f(\color{royalblue}{x+h})=\frac{1}{\color{royalblue}{x+h}+3}\]Understand that part? :)

OpenStudy (anonymous):

sorry i will ba back in 30 mins so sorry

zepdrix (zepdrix):

lol c: k

OpenStudy (anonymous):

Hello are you there?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

\[f'(x)=\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\\ \color{red}{f'(x)=\lim_{h\to 0}\frac{\frac{1}{x+h+3}-\frac{1}{x+3}}{h}}\\ f'(x)=\lim_{h\to 0}\frac{\frac{x+3}{(x+h+3)(x+3)}-\frac{x+h+3}{(x+3)(x+h+3)}}{h}\\ f'(x)=\lim_{h\to 0}\frac{\frac{x+3-x-h-3}{(x+h+3)(x+3)}}{h}\\ f'(x)=\lim_{h\to 0}\frac{\frac{x+3-x-h-3}{(x+h+3)(x+3)}}{h}\] Does that help? The red part is what @zepdrix was talking about.

OpenStudy (anonymous):

no im confused from the start, why its 1/x+h+3?

OpenStudy (anonymous):

Like @zepdrix said, \[f(x)=\frac{1}{x+3}, \text{ so}\\ f(x+h)=\frac{1}{(x+h)+3}\]

OpenStudy (anonymous):

Oh i see :)

OpenStudy (anonymous):

thank you! can i ask more question?

OpenStudy (anonymous):

Start another thread. It's getting messy with all these replies.

OpenStudy (anonymous):

ok, sure i will make a new state